OpenStudy (love_ranaa):

Medal & Fan @PizzaLover123 @jeffyblood

OpenStudy (anonymous):

Whats the question xD

OpenStudy (anonymous):

The question? :)

OpenStudy (anonymous):

Please show us the question. It makes it easier to solve. :) lol

OpenStudy (anonymous):

@DIDITHEDAD @Dethmew Ok she heard us LOL.

OpenStudy (love_ranaa):

Sorry..I closed everything

OpenStudy (love_ranaa):

How are you supposed to do that ...

OpenStudy (anonymous):

Do you have a preference?

OpenStudy (anonymous):

Oh, too be honest I have no idea XP. Maybe someone who knows what he is talking about should help you XP

OpenStudy (anonymous):

I have no idea ALSO LOL @iGreen

OpenStudy (igreen):

Use the distance formula.

OpenStudy (anonymous):

Usually its a subtracting formula.

OpenStudy (igreen):

\(\sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (igreen):

Plug in the two points and simplify.

OpenStudy (love_ranaa):

Oh, okay..Thanks !

OpenStudy (igreen):

Happy to Help \(\Huge\ddot\smile\)

OpenStudy (love_ranaa):

What comes first..5,6 or 1,1..

OpenStudy (anonymous):

5 - 1 = 4 6 - 1 = 5 5 + 4 =

OpenStudy (anonymous):

9? @iGreen

OpenStudy (igreen):

(1, 1), (5, 6) x1 y1 x2 y2

OpenStudy (igreen):

@Love_Ranaa

OpenStudy (igreen):

\(\sqrt{(5-1)^2 + (6-1)^2}\)

OpenStudy (love_ranaa):

Okay, I got B. 6.4

OpenStudy (igreen):

Correct.

OpenStudy (igreen):

5 - 1 = 4 6 - 1 = 5 4^2 = 16 5^2 = 25 16 + 25 = 41 sqrt of 41 is approximately 6.4

OpenStudy (love_ranaa):

Okay..Thanks you soooo much @iGreen And your tagging helped too @PizzaLover123

OpenStudy (igreen):

No problem, Happy to Help \(\Huge\ddot\smile\)

OpenStudy (anonymous):

Sorry I couldn't help more but LOL Thanks :D And @iGreen thanks for answering when We tag you.